Could you please provide the detailed solution with full explanation for the final decision.

The corporation, which is currently operating at full capacity, has sales of $47,000, current assets of $5,100, current liabilities of $6,200, net fixed assets of $51,500, and a 3.1 percent profit margin. The firm has no long-term debt and does not plan on acquiring any. The firm does not pay any dividends. Sales are expected to increase by 3 percent next year. If all assets, short-term liabilities, and costs vary directly with sales, how much additional equity financing is required for next year?

A) $ 0

B) $ 500

C) $ 967

D) $ 1698

E) $ 1512
Answer #1

1)Next year sales : 47000 (1+.03) = 48410

Next year profit margin = 48410*.031 = 1500.71 [This will be an addition to retained earning since there is no dividend payout]

2)current year :

Total asset = 5100+ 51500= 56600

Total liabilities =6200

Total equity = 56600-6200= 50400

Next year :

Total asset : 56600(1+.03)= 58298

Total liabilities = 6200(1+.03) = 6386

Equity = 58298-6386 = 51912

Equity = Equity last year + addition to retained earning + AFN

51912 =50400+1500.71+ AFN

AFN = 51912-50400-1500.71

      = 11.29    (almost nearest to 0 since 11.29 is minimal amount]

Correct option is "A" - 0